OpenStudy (anonymous):

A tank of water holding 528 kL of water begins to empty at a rate of 10 kL/min. At the same time, another tank, which is empty, begins to fill at a rate of 14 kL/min. Let k represent the number of kiloliters of water and let t represent time in minutes. The system models this situation. How long will it take for the tanks to hold equal amounts of water, and how much water will that be? k = 528 – 10t k = 14t It will take minutes__ for both tanks to hold equal amounts of water. They will each hold__ kiloliters. @iGreen

OpenStudy (igreen):

Just solve 528 - 10t = 14t: 528 - 10t = 14t Add 10t to both sides: 528 = 24t Divide 24 to both sides: What's 528 / 24?

OpenStudy (anonymous):

22. 528/24=22

OpenStudy (igreen):

@Rosabearof10 I was asking @NunChuckBoy3101

OpenStudy (anonymous):

dont give me the answer i am trying to get help not the answer

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

i need a challenge so i can get it

OpenStudy (igreen):

What's 528 / 24? @NunChuckBoy3101

OpenStudy (anonymous):

but its ok @iGreen i already know the answer now so whats next

OpenStudy (anonymous):

22

OpenStudy (anonymous):

I was just trying to help. Sorry, i will leave u alone.

OpenStudy (igreen):

Okay. You put 22 in the first blank. That's how much time it takes for both equations to reach the same amount. Next we plug 22 into the second equation: k = 14t k = 14(22) What's 14 * 22? @NunChuckBoy3101

OpenStudy (igreen):

@Rosabearof10 I know you're trying to help, but you're not supposed to give out answers.

OpenStudy (anonymous):

goodbye. -just helping-

OpenStudy (anonymous):

no its ok i know just explain the problem then make them work it out don't just give them the answer

OpenStudy (igreen):

^^

OpenStudy (anonymous):

But is this what open study is for?

OpenStudy (anonymous):

14*22=308

OpenStudy (igreen):

Yes! So that's what we put in our 2nd blank..it's the amount of water both equations hold.

OpenStudy (anonymous):

there are free tutors on here that help you get what subject that you need help in

OpenStudy (anonymous):

what do you mean @iGreen

OpenStudy (anonymous):

@iGreen what do you mean

OpenStudy (igreen):

Huh..? First blank = 22 2nd blank = 308 That means: It will take minutes 22 for both tanks to hold equal amounts of water. They will each hold 308 kiloliters.

OpenStudy (anonymous):

oh ok thanks